# Detecting Measurement-Induced Entanglement Transitions With Unitary Mirror Circuits

### v1

@@ Abstract
Monitored random circuits, consisting of alternating layers of random entangling unitaries and local measurement operations on a randomly chosen subsets $p$ of all qubits, have recently come into focus as a probe into the interplay of unitary and non-unitary evolution, quantum error correcting codes, and more [Skinner2019, Szyniszewski2020, Choi2020, Gullans2020, Gopalakrishnan2021, Ippoliti2021, Bao2022]. These circuits have been shown to exhibit a phase transition between long range, volume-law entanglement between the qubits at $p<p_{\rm c}$ and a local, area-law entanglement phase at $p>p_{\rm c}$, for some critical measurement density $p_{\rm c}$ [Skinner2019]. Significantly, this is not a property of the entanglement of the statistical ensemble of states; it is only seen if one averages the entanglement of each individual outcome, characterized by the circuit and the record of measurement results [Friedman2023]. This makes it difficult to experimentally access these properties, as generally the number of repetitions required to recreate same state twice grows exponentially with the number of measurements.
((a))
((b))
((c))
@@
Figure 1: Outline of our proposed scheme. LABEL:sub@fig:sketchMIPTcirc A randomly monitored circuit is run on a quantum device generating a state $\lvert\psi\rangle$. LABEL:sub@fig:sketchMPSprocess The gate parameters and measurement results are fed into a classical processor, generating a deterministic circuit where measurements have been replaced with projection operators (marked as diamonds). We classically calculate an MPS $\lvert\psi_{D}\rangle$ approximating the output of this circuit, and from it a “unitary mirror” circuit that generates it, $\mathcal{C}_{D}\sim\lvert\psi_{D}\rangle\langle 0^{\otimes N}\rvert$. LABEL:sub@fig:sketchMPScirc Finally, we apply the inverted mirror $\mathcal{C}_{D}^{\dagger}$ to $\lvert\psi\rangle$ and measure the probability of returning to the initially state, experimentally measuring the overlap $\mathopen{}\mathclose{{}\left\lvert\langle\psi_{D}|\psi\rangle}\right\rvert^{2}$ averaged over all runs.
@@
A variety of lateral approaches have been suggested to overcome this difficulty [Gullans2020a, Lee2022, Noel2022, Hoke2023, Dehghani2023, Garratt2023, Li2023a]. Invariably, they rely on some classical calculation that attempts to recreate the quantum dynamics of the circuit. In principle, for any instance of a monitored circuit, consisting of an initial state, a set of unitary operations and measurement points, and a record of measurement results, the final state of the system is deterministic, and could be obtained by a classical computer. In practice, this generically requires calculating the unitary matrix of a many-qubit operation, and thus one has simply shifted the exponential requirement from the experimental regime into the computational. To avoid this overhead, experimental approaches generally use an approximation of the state to extract the entropy, introducing errors or uncertainty to the result.
@@
We consider a chain of $N$ qubits and an $L$-layer circuit $\mathcal{C}$. The circuit is composed of a set of gates, $\mathbf{U}=\{\hat{U}_{\ell,2q}|1\leq\ell\leq L,1\leq q\leq\lfloor N/2\rfloor\}$, sampled out of some set of two-qubit unitaries $\hat{U}_{\ell,q}\in\mathcal{U}$, and measurement points, =$\mathcal{M}=\{M_{\ell}|1\leq\ell\leq L\}$, where $M_{\ell}\subset\{1,\dotsc,N\}$, $\mathopen{}\mathclose{{}\left\lvert M_{\ell}}\right\rvert=pN$ for some measurement rate $p$. At each odd (even) layer $\ell$, we apply $\hat{U}_{\ell,q}$ between each even qubit $2q$ and its preceding (following) neighbor, $2q+\mathopen{}\mathclose{{}\left({-1}}\right)^{\ell}$. After this, we measure each qubit $q\in M_{i}$ recording the results into an $L\times pN$ matrix $\mathbf{m}$. The circuit generates a final state $\lvert\psi\mathopen{}\mathclose{{}\left[{\mathbf{U},\mathbf{M},\mathbf{m}}}\right]\rangle$.
@@
To probe entropy properties, we extend the circuit as follows. We run $\mathcal{C}$ as given above, generate $\lvert\psi\mathopen{}\mathclose{{}\left[{\mathbf{U},\mathbf{M},\mathbf{m}}}\right]\rangle$, and keep it in quantum memory. Then $\mathbf{U}$, $\mathbf{M}$, and $\mathbf{m}$ are sent to a classical computer, which calculates an approximate final state $\lvert\psi_{D}\mathopen{}\mathclose{{}\left[{\mathbf{U},\mathbf{M},\mathbf{m}}}\right]\rangle$ using a matrix product state (MPS) decomposition with bond dimension $D$ [Cirac2021]. We convert it [Schon2005, Ran2020] into a “unitary mirror” $\mathcal{C}_{D}$, a unitary-only circuit that when applied to the same initial state, generates $\lvert\psi_{D}\rangle$. The quantum device then applies the inverse of the mirror to the state. Finally, we measure the probability of returning to the initial state. By repeating this process, we measure the mirror fidelity
$F_{D}\mathopen{}\mathclose{{}\left[{\mathbf{U},\mathbf{M},\mathbf{m}}}\right]={\mathopen{}\mathclose{{}\left\lvert\langle\psi_{D}\mathopen{}\mathclose{{}\left[{\mathbf{U},\mathbf{M},\mathbf{m}}}\right]|\psi\mathopen{}\mathclose{{}\left[{\mathbf{U},\mathbf{M},\mathbf{m}}}\right]\rangle}\right\rvert^{2}},$ (1)
averaged over all possible $\mathbf{m}$.
@@
Next, consider the Von Neuman entanglement entropy on the subset, $S^{(n)}=-\sum_{i}\mu^{(n)}_{i}\log\mu^{(n)}_{i}$. For a given Schmidt error, the entropy can be maximized by taking $\mu_{i\leq D}=(1-\epsilon^{(n)}_{D})/D$, $\mu_{i>D}=\epsilon^{(n)}_{D}/(2^{\bar{n}}-D)$. Thus, given $\epsilon^{(n)}_{D}$, the entanglement entropy of the subset $i=1\dotsc n$ is bounded by [supplemental]
@@
Combining the two bounds, we arrive at our main result. We find that given a state approximated by an MPS of bond dimension $D$ with mirror fidelity $F$, the entanglement entropy for half the chain is bounded by
@@
Figure 3: The upper bound on the value of the volume-law coefficient as derived from the MPS protocol described in the text. Shown for circuits using (a-d) randomly chosen two-qubit Clifford gates and (e-h) Haar-random two-qubit gates as entanglers. The dotted gray lines denote the critical point $p_{\rm c}$ [Skinner2019, Gullans2020]. (a-c), (e-g) Each subplot corresponds to the noted chain length $N$, and each curve corresponds to a different bond dimension $D$. We observe the $\log D$ behavior implied by Eq. 5. (d), (h) We extract an overall curve for each $N$ by taking the minimum values of $F$ over $D$. We can see how the bound on the volume law decreases as we increase $N$; this takes the form of the $1/N$ dependence shown in Eq. 6.
@@
We numerically explore this bound in Fig. 3 for two sets of gates $\mathcal{U}$ [Johansson2013, Roberts2019, Gidney2021]. For two-qubit gates sampled from the set of Haar-random gates, we consider small chains, up to $N=16$. For $\mathcal{U}$ consisting of the set of Clifford gates, we consider up to $N=60$. In both cases we consider various bond dimensions up to $D=N$. We observe, in both cases, behavior consistent with the real bound, shown as a black line.
@@
Ref. Verstraete2006 has shown that there exists some MPS approximation that has $1-\sum_{n=1}^{N-1}\epsilon^{(n)}_{D}\leq\sqrt{F}$. However, this is a very loose bound, as evidenced by the left hand side quickly becoming negative for large values of $N$, and so is the resulting bound on the entropy, $S^{(N/2)}\gtrsim\mathopen{}\mathclose{{}\left({\log D}}\right)/N$. Beyond this, the bound on the fidelity only applies to the best possible MPS used to approximate the state, which we do not know how to construct from the generating circuit in polynomial time.
Instead, let us consider the evolution of the approximate wavefunction along the circuit. We obtain $\lvert\psi_{D}\rangle$ by a simple iterative process: we begin with the initial state used in the circuit, $\lvert\psi_{D}^{0}\rangle=\lvert 0\rangle^{\otimes N}$. Then, given an approximate state $\lvert\psi_{D}^{\ell-1}\rangle$, we generate $\lvert\psi_{D}^{\ell}\rangle$ by applying the gates and projection operators of layer $\ell$, and then truncating the state back to bond dimension $D$. We can obtain the truncation error of these individual steps as we perform the calculation, which will be smaller than the total Schmidt error. Thus, we can substitute it into Eq. 7 to obtain a lower bound on the entropy of the real state.
@@
Figure 4: The lower bound on the value of the volume-law coefficient as derived from the MPS protocol described in the text. Shown for circuits using (a-d) randomly chosen two-qubit Clifford gates and (e-h) Haar-random two-qubit gates as entanglers. The dotted gray lines denote the critical point $p_{\rm c}$. Each subplot corresponds to the noted chain length $N$, and each curve corresponds to a different bond dimension $D$.
Notably, this lower bound estimate makes no use of $F$; it relies entirely on feedback from the numerical estimate of the MPS. This obviously reduces the quantum overhead required, but it also reduces the classical computation somewhat, as a smaller light-cone style circuit can be calculated to extract only ${\epsilon^{(N/2)}_{D}}$. Still, this bound is logarithmic in $D$ and so the resource requirements scale unfavorably. However, as mentioned above, the upper bound, which has polynomial requirements, appears to be tight and may prove sufficient to extract the critical point $p_{\rm c}$.
@@
We consider the computational resources required by our proposed protocol, both classical and quantum. The quantum resources required are outlines in Fig. 1. These are polynomial in $N$ by construction: the application of the monitored random circuit requires a depth of $L\propto N$ layers, with each layer having $N/2$ concurrent two-qubit gates and $pN$ concurrent measurements. The inverted circuit can be straightforwardly decomposed into $N$ unitary $\log_{2}D$-qubit operations [Schon2005, Ran2020], which can be further decomposed into $O(D^{2})$ two-qubit gates each [Shende2006, Krol2022].
@@
The probing of the lower bound estimate, which relies only on the construction of the MPS, scales less favorably, as the bound in Eq. 7 grows as $\log D$. Using the same $\varepsilon$-test to find a minimum coefficient on the log portion of the entropy, $S^{(N/2)}>\varepsilon\log N/2$, the bond dimension would have to be scaled as $D\sim N^{1/\varepsilon}$, which is polynomial but may be quite large. It is possible that such calculations can be done more effectively, as recent work has often shown that MPS methods for specific problems can be greatly improved [Ayral2023, Tindall2023].
@@
Finally, we note that the resources needed to produce a quantum-advantage result here are not large. The quantum resources, in particular, are simply linear in $N$, both in the number of qubits required and the depth of the algorithm. The biggest roadblock to implementing such algorithms is the need to implement quantum-classical communication mid-circuit, and go beyond simple if statements in the quantum code. Such capabilities would open the door to many more hybrid algorithms.
@@ I Maximum fidelity of an MPS
$\begin{gathered}\lambda_{i}\geq\lambda_{i+1},\qquad\sum_{i=1}^{2^{\bar{n}}}\mu_{i}=1,\\ \lvert\alpha_{i}\rangle=\sum_{j}U_{ij}\lvert A_{j}\rangle,\qquad\lvert\beta_{i}\rangle=\sum_{j}V_{ij}\lvert B_{j}\rangle,\end{gathered}$
@@ II Relations between $\epsilon^{(n)}_{D}$ and the entropy
$S^{(n)}_{\alpha}=\frac{1}{1-\alpha}\log\mathopen{}\mathclose{{}\left[{\operatorname{Tr}\rho_{A}^{\alpha}}}\right]=\frac{1}{1-\alpha}\log\mathopen{}\mathclose{{}\left[{\sum_{i=1}^{2\bar{n}}\mu_{i}^{\alpha}}}\right],$ (S5)

### v2

@@ Abstract
Monitored random circuits, consisting of alternating layers of random entangling unitaries and local measurement operations on a randomly chosen subsets $p$ of all qubits, have recently come into focus as a probe into the interplay of unitary and non-unitary evolution, quantum error correcting codes, and more [Skinner2019, Szyniszewski2020, Choi2020, Gullans2020, Gopalakrishnan2021, Ippoliti2021, Bao2024]. These circuits have been shown to exhibit a phase transition between long range, volume-law entanglement between the qubits at $p<p_{\rm c}$ and a local, area-law entanglement phase at $p>p_{\rm c}$, for some critical measurement density $p_{\rm c}$ [Zabalo2020]. Significantly, this is not a property of the entanglement of the statistical ensemble of states; it is only seen if one averages the entanglement of each individual outcome, characterized by the circuit and the record of measurement results [Friedman2023]. This makes it difficult to experimentally access these properties, as generally the number of repetitions required to recreate same state twice grows exponentially with the number of measurements.
@@
Figure 1: Outline of our proposed scheme. (a) A randomly monitored circuit is run on a quantum device generating a state $\lvert\psi\rangle$. (b) The gate parameters and measurement results are fed into a classical processor, generating a deterministic circuit where measurements have been replaced with projection operators (marked as diamonds). We classically calculate an MPS $\lvert\psi_{D}\rangle$ approximating the output of this circuit, and from it a “unitary mirror” circuit that generates it, $\mathcal{C}_{D}\lvert 0^{\otimes N}\rangle=\lvert\psi_{D}\rangle$. (c) Finally, we apply the inverted mirror $\mathcal{C}_{D}^{\dagger}$ to $\lvert\psi\rangle$ and measure all qubits. (d) The probability of finding the all-zero state is proportional to the overlap $\mathopen{}\mathclose{{}\left\lvert\langle\psi_{D}|\psi\rangle}\right\rvert^{2}$.
@@
A variety of lateral approaches have been suggested to overcome this difficulty. One tack has been to apply a form of steering, or preselection, to direct the system into a single state in the entangled regime [Buchhold2022, Iadecola2023, Piroli2023, Sierant2023]. While this allows for experimental implementation, the dynamics and universality regime generally diverge from those of undirected MIPT formulations. Techniques looking to probe the undirected form usually rely on some classical calculation that attempts to recreate the quantum dynamics of the circuit, and correlate the final measurement with that calculation [Gullans2020a, Lee2022, Noel2022, Hoke2023, Dehghani2023, Garratt2023, Li2023a]. In principle, for any instance of a monitored circuit, consisting of an initial state, a set of unitary operations and measurement points, and a record of measurement results, the final state of the system is deterministic, and could be obtained by a classical computer. In practice, this generically requires calculating the unitary matrix of a many-qubit operation, and thus one has simply shifted the exponential requirement from the experimental regime into the computational. To avoid this overhead, experimental approaches generally use an approximation of the state to extract the entropy, introducing errors or uncertainty to the result.
@@
We consider a chain of $N$ qubits and an $L$-layer circuit $\mathcal{C}$. The circuit is composed of a set of gates, $\mathbf{U}=\{\hat{U}_{\ell,2q}|1\leq\ell\leq L,1\leq q\leq\lfloor N/2\rfloor\}$, sampled out of some set of two-qubit unitaries $\hat{U}_{\ell,q}\in\mathcal{U}$, and measurement points, $\mathbf{M}=\{M_{\ell}|1\leq\ell\leq L\}$, where $M_{\ell}\subset\{1,\dotsc,N\}$, $\mathopen{}\mathclose{{}\left\lvert M_{\ell}}\right\rvert=pN$ for some measurement rate $p$. At each odd (even) layer $\ell$, we apply $\hat{U}_{\ell,q}$ between each even qubit $2q$ and its preceding (following) neighbor, $2q+\mathopen{}\mathclose{{}\left({-1}}\right)^{\ell}$. After this, we measure each qubit $q\in M_{\ell}$ recording the results into an $L\times pN$ matrix $\mathbf{m}$. The circuit generates a final state $\lvert\psi\mathopen{}\mathclose{{}\left[{\mathbf{U},\mathbf{M},\mathbf{m}}}\right]\rangle$.
To probe entropy properties, we extend the circuit as follows. We run $\mathcal{C}$ as given above, generate $\lvert\psi\rangle$, and keep it in quantum memory. Then $\mathbf{U}$, $\mathbf{M}$, and $\mathbf{m}$ are sent to a classical computer, which calculates an approximate final state $\lvert\psi_{D}\rangle$ using a matrix product state (MPS) decomposition with bond dimension $D$ [Cirac2021]. We convert it [Schon2005, Ran2020, supplemental] into a “unitary mirror”, a unitary-only circuit that when applied to an all-zero states, generates $\mathcal{C}_{D}\lvert 0^{\otimes N}\rangle=\lvert\psi_{D}\rangle$. The quantum device then applies the inverse of the mirror to the state, to generate $\lvert\phi_{D}\rangle=\mathcal{C}_{D}^{\dagger}\lvert\psi\rangle$. Finally, we measure the probability of finding all qubits in the zero state, to obtain the overlap
$\begin{split}F_{D}\mathopen{}\mathclose{{}\left[{\mathbf{U},\mathbf{M},\mathbf{m}}}\right]&=\mathopen{}\mathclose{{}\left\lvert\langle 0^{\otimes N}|\phi_{D}\mathopen{}\mathclose{{}\left[{\mathbf{U},\mathbf{M},\mathbf{m}}}\right]\rangle}\right\rvert^{2}\\ &=\mathopen{}\mathclose{{}\left\lvert\langle 0^{\otimes N}\rvert\mathcal{C}_{D}^{\dagger}\mathopen{}\mathclose{{}\left[{\mathbf{U},\mathbf{M},\mathbf{m}}}\right]\lvert\psi\mathopen{}\mathclose{{}\left[{\mathbf{U},\mathbf{M},\mathbf{m}}}\right]\rangle}\right\rvert^{2}\\ &=\mathopen{}\mathclose{{}\left\lvert\langle\psi_{D}\mathopen{}\mathclose{{}\left[{\mathbf{U},\mathbf{M},\mathbf{m}}}\right]|\psi\mathopen{}\mathclose{{}\left[{\mathbf{U},\mathbf{M},\mathbf{m}}}\right]\rangle}\right\rvert^{2}.\end{split}$ (1)
By repeating this process, we measure the mirror fidelity averaged over all possible outcomes $\mathbf{m}$.
@@
Next, consider the Von Neuman entanglement entropy on the subset, $S^{(n)}=-\sum_{i}\mu^{(n)}_{i}\log\mu^{(n)}_{i}$. For a given Schmidt error, the entropy can be maximized by taking $\mu_{i\leq D}=(1-\epsilon^{(n)}_{D})/D$, $\mu_{i>D}=\epsilon^{(n)}_{D}/(2^{\bar{n}}-D)$. Thus, given $\epsilon^{(n)}_{D}$, the entanglement entropy of the subset $i=1\dotsc n$ is bounded by
@@
Figure 3: The average overlap $F$ of the unitary mirror generated by the MPS for various random circuits. Shown for circuits using (a-c) randomly chosen two-qubit Clifford gates and (d-f) Haar-random two-qubit gates as entanglers. The dotted gray lines denote the critical point $p_{\rm c}$ [Sierant2022]. Each subplot corresponds to the noted chain length $N$, and each curve corresponds to a different bond dimension $D$. For the Clifford circuits, we see a clear inflection point at the critical density; for the Haar circuits, we see signs of a similar phenomenon.
Figure 4: The upper bound on the value of the volume-law coefficient, derived from overlaps shown in Fig. 3, as described in the text. Shown for circuits using (a-c,g) randomly chosen two-qubit Clifford gates and (d-f,h) Haar-random two-qubit gates as entanglers. The dotted gray lines denote the critical point $p_{\rm c}$ [Sierant2022]. (a-f) Each subplot corresponds to the noted chain length $N$, and each curve corresponds to a different bond dimension $D$, corresponding to the same plot in Fig. 3. We observe the $\log D$ behavior implied by Eq. 5. (g-h) We extract an overall curve for each $N$ by taking the minimum value of $S$ over $D$. We can see how the bound on the volume law decreases as we increase $N$; this takes the form of the $1/N$ dependence shown in Eq. 6.
Combining the two bounds [supplemental], we arrive at our main result. We find that given a state approximated by an MPS of bond dimension $D$ with mirror fidelity $F$, the entanglement entropy for half the chain is bounded by
@@
We numerically explore this bound in Figs. 3 and 4 for two sets of gates $\mathcal{U}$ [Johansson2013, Roberts2019, Gidney2021]. For two-qubit gates sampled from the set of Haar-random gates, we consider small chains, up to $N=16$. For $\mathcal{U}$ consisting of the set of Clifford gates, we consider up to $N=60$. In both cases we consider various bond dimensions up to $D=N$. We observe, in both cases, behavior consistent with the real bound, shown as a black line.
@@
Unfortunately, we have found no no easy way to bound $\epsilon^{(n)}_{D}$ from below with the fidelity $F$. Ref. Verstraete2006 has shown that there exists a specific MPS construction that has $1-\sum_{n=1}^{N-1}\epsilon^{(n)}_{D}\leq\sqrt{F}$, but the time to calculate it from the given circuit scales exponentially with $N$. In addition, this is a very loose bound, as evidenced by the left hand side quickly becoming negative for large values of $N$.
Instead, let us consider the evolution of the approximate wavefunction along the circuit. We obtain $\lvert\psi_{D}\rangle$ by a simple iterative process [supplemental]: we begin with the initial state used in the circuit, $\lvert\psi_{D}^{0}\rangle=\lvert 0\rangle^{\otimes N}$. Then, given an approximate state $\lvert\psi_{D}^{\ell-1}\rangle$, we generate $\lvert\psi_{D}^{\ell}\rangle$ by applying the gates and projection operators of layer $\ell$, and then truncating the state back to bond dimension $D$. We can obtain the truncation error of these individual steps as we perform the calculation, which will be smaller than the total Schmidt error. Thus, we can substitute it into Eq. 7 to obtain a lower bound on the entropy of the real state.
@@
Notably, this lower bound estimate makes no use of $F$; it relies entirely on feedback from the numerical estimate of the MPS. This obviously reduces the quantum overhead required, but it also reduces the classical computation somewhat, as a smaller light-cone style circuit can be calculated to extract only ${\epsilon^{(N/2)}_{D}}$. Still, this bound is logarithmic in $D$ and so the resource requirements scale unfavorably. However, as mentioned above, the upper bound, which has polynomial requirements, appears to be tight and may prove sufficient to extract the critical point $p_{\rm c}$. Alternately, one may combine the upper bound from $F$ with lower bounds estimated by some other means [Garratt2023, McGinley2023].
@@
We consider the computational resources required by our proposed protocol, both classical and quantum. The quantum resources required are outlined in Fig. 1. These are polynomial in $N$ by construction: the application of the monitored random circuit requires a depth of $L\propto N$ layers, with each layer having $N/2$ concurrent two-qubit gates and $pN$ concurrent measurements. The inverted circuit can be straightforwardly decomposed into $N$ unitary $\log_{2}D$-qubit operations [supplemental, Schon2005, Ran2020], which can be further decomposed into $O(D^{2})$ two-qubit gates each [Shende2006, Krol2022] or $\tilde{O}(\sqrt{D})$ gates and $O\mathopen{}\mathclose{{}\left({D}}\right)$ ancillas [Rosenthal2023].
@@
Throughout the paper we have focused on the common case of a one-dimensional chain. In fact, the derivation described in Eqs. 3 to 6 generalizes to higher dimensional systems. The number of terms in the Schmidt decomposition of the tensor network simply goes from $D\to D^{\mathopen{}\mathclose{{}\left\lvert A}\right\rvert}$, where $\mathopen{}\mathclose{{}\left\lvert A}\right\rvert$ is the number of qubits in the boundary of the subset, and the same bounds follow. However, it may not be straightforward to generate an appropriate high-fidelity MPS in polynomial time and convert it into a unitary mirror circuit with a polynomial number of gates. These may still be possible with specific forms of the tensor network [Zaletel2020].
@@
Alternatively, one can imagine an version of the scheme where the mirror circuit is applied directly to a set of auxiliary qubits, generating the state $\lvert\psi\rangle\otimes\lvert\psi_{D}\rangle$. This could be used as a resource in any number of generalized schemes to probe the monitored circuit ensemble [McGinley2023].
@@
Finally, we note that the resources needed to produce a quantum-advantage result here are not large. The quantum resources required for the unitary mirror, in particular, may be as little as $D$ ancilla qubits and a depth of $O(\sqrt{D})$ gates [supplemental, Rosenthal2023]. Thus, they are likely to be smaller than those required for the monitored circuit itself. The other roadblock to implementing such algorithms is the need to implement quantum-classical communication mid-circuit, and go beyond simple if statements in the quantum code. Such capabilities would open the door to many more hybrid algorithms.
@@ I Maximum fidelity of an MPS
$\begin{gathered}\lambda_{i}\geq\lambda_{i+1},\qquad\sum_{i=1}^{2^{\bar{n}}}\lambda_{i}=1,\\ \lvert\alpha_{i}\rangle=\sum_{j}U_{ij}\lvert A_{j}\rangle,\qquad\lvert\beta_{i}\rangle=\sum_{j}V_{ij}\lvert B_{j}\rangle,\end{gathered}$
@@ II Relations between $\epsilon^{(n)}_{D}$ and the entropy
$S^{(n)}_{\alpha}=\frac{1}{1-\alpha}\log\mathopen{}\mathclose{{}\left[{\operatorname{Tr}\rho_{A}^{\alpha}}}\right]=\frac{1}{1-\alpha}\log\mathopen{}\mathclose{{}\left[{\sum_{i=1}^{2^{\bar{n}}}\mu_{i}^{\alpha}}}\right],$ (S5)